OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (-4, 0) and a directrix at x = 4.

OpenStudy (amistre64):

all points that are the same distance from the focus and the directrix .... equate the distances

OpenStudy (anonymous):

the distance from the focus to the directrix is 8

OpenStudy (amistre64):

\[d_f^2=(x+4)^2+(y)^2\] \[d_d^2=(x-4)^2+(y-y)^2\]

OpenStudy (amistre64):

\[(x+4)^2 + y^2=(x-4)^2\] \[x^2+8x+16 + y^2=x^2-8x+16\] \[\cancel{x^2}+8x\cancel{+16} + y^2=\cancel{x^2}-8x\cancel{+16}\] \[8x+y^2=-8x\] etc....

OpenStudy (anonymous):

would this be the final answer

OpenStudy (amistre64):

of course not ... hence the 'etc...'

OpenStudy (anonymous):

ok so I would have to divide by -8 on both sides

OpenStudy (amistre64):

you could if you wanted to ... work out the steps and let me see what you come to

OpenStudy (amistre64):

are we getting stuck in another site crash?

OpenStudy (anonymous):

8x+y^2=-8x y^2=-1x

OpenStudy (anonymous):

idk

OpenStudy (amistre64):

your thought was fine, your process if off

OpenStudy (amistre64):

8x +y^2 = -8x /-8 /-8 /-8 --------------- -x - y^2/8 = x might i suggest adding and x to both sides?

OpenStudy (anonymous):

y^2/8=x^2

OpenStudy (amistre64):

a different approach would be to get the xs all to one side first 8x+y^2=-8x -8x -8x -------------- y^2 = -16x then divide, just less steps to take

OpenStudy (amistre64):

-x - y^2/8 = x +x +x -------------- -y^2/8 = 2x adding is not multiplication so no x^2

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

i take it you got it from here? good luck